# Chiara Amorino, Vytautė Pilipauskaitė

### v2

@@ 3 Estimators and main results
We address the condition $N\Delta_{n}\rightarrow 0$, which is crucial for the asymptotic normality in Theorem 3.2. This requirement on $\Delta_{n}$ has been extensively studied in the context of classical SDEs with discrete observations over a time horizon $T:=n\Delta_{n}\to\infty$. In [29], the condition was $n\Delta_{n}^{2}\rightarrow 0$, which was later improved to $n\Delta_{n}^{3}\rightarrow 0$ in [85]. Kessler [54] further relaxed it to $n\Delta_{n}^{p}\rightarrow 0$ using a Gaussian approximation. Similar developments are found in SDEs with jumps [5, 6, 36, 78]. Reinforcing the parallel between $N$ and $T$, this same condition appears in [4], where potential improvements involving higher-order approximations were discussed. For us, the condition $N\Delta_{n}\rightarrow 0$ is primarily needed to approximate the contrast derivative with martingale increments, as in classical SDEs, but also to manage particle correlations. While higher-order approximations could relax the condition in the analysis of the contrast derivative, they do not fully resolve the issue of particle correlation. In our case, the complexity is heightened by the dependency between $X$ and $\tilde{X}$ in partial observations, making further relaxation challenging. We leave this exploration for future work.
Remark 3.8.
We emphasize that our primary goal is to initiate the development of statistical inference methods for interacting particle systems, recognizing that there is still a long way to go to reach the level of maturity seen in the non-interacting case. In this manuscript, we adopt a locally degenerate approximation for constructing the contrast function, relying on the Euler–Maruyama discretization of the true, intractable process, whose covariance matrix is locally degenerate.
An alternative approach, used in the non-interacting setting (see e.g., [26, 37, 47, 63, 71]), is to consider a locally non-degenerate approximation, where a joint approximate transition density—different from the EM scheme—with a non-degenerate covariance is employed in the contrast function. This strategy enables inference in more general hypoelliptic models. However, the asymptotic analysis in such a framework is significantly more delicate and requires more sophisticated tools.
For this reason, and to keep the scope of the present manuscript focused and manageable, we have opted to proceed with the locally degenerate approximation. The extension to the locally non-degenerate case for IPS is a promising direction. Interestingly, the very recent contribution [50], which appeared during the revision of our manuscript, develops precisely along these lines. We see this as further evidence of both the timeliness and impact of our work, which has already stimulated extensions toward broader classes of degenerate IPS.
@@ 4 On the Hormander’s condition for hypoellipticity
First, let us introduce the empirical projection function. Given a real-valued function $\phi$ over $\mathcal{P}_{2}(\mathbb{R}^{l})$ for some $l\in\mathbb{N}$, for every $N\in\mathbb{N}$ we define $\phi^{N}:\mathbb{R}^{l\times N}\rightarrow\mathbb{R}$ by:
$\phi^{N}(z_{1},\ldots,z_{N}):=\phi\left(\frac{1}{N}\sum_{j=1}^{N}\delta_{z_{j}}\right).$
Additionally, let us denote the projection onto the $i$-th component by $\pi^{(i)}$, so that $\pi^{(i)}:\mathbb{R}^{2N}\rightarrow\mathbb{R}^{2}$ is defined as:
$\pi^{(i)}\big{(}(x_{1},y_{1}),\ldots,(x_{N},y_{N})\big{)}=(x_{i},y_{i}).$
Then, for $j=1,2$, we can express:
$b_{j}(Z_{t}^{i},\Pi_{t}^{N})=b_{j}\left(Z_{t}^{i},\frac{1}{N}\sum_{j=1}^{N}\delta_{Z_{t}^{j}}\right)=b_{j}\big{(}\pi^{(i)}(Z_{t}),I_{d}^{N}(Z_{t})\big{)}=:b_{j}^{(i)}(Z_{t}),$
where we denote the identity function by $I_{d}$ and the vector $(Z_{t}^{1},\ldots,Z_{t}^{N})$ by $Z_{t}$. Similarly, $a(Z_{t}^{i},\Pi_{t}^{N})$ can be expressed as $a^{(i)}(Z_{t})$.
@@
$dZ_{t}=B(Z_{t})dt+A(Z_{t})dB_{t},$ (10)
where $B_{t}$ is a $2N$-dimensional Brownian motion. Here,
$B(Z_{t})=(b_{1}^{(1)}(Z_{t}),b_{2}^{(1)}(Z_{t}),...,b_{1}^{(N)}(Z_{t}),b_{2}^{(N)}(Z_{t}))\in\mathbb{R}^{2N},$
and $A=(A_{1}|A_{2}|...|A_{2N})$ is a $2N\times 2N$ matrix with $A_{2k+1}=\mathbf{0}\in\mathbb{R}^{2N}$ for $k\in\{0,...,N-1\}$, while for $k\in\{1,...,N\}$, $A_{2k}=(0,...,a^{(k)}(Z_{t}),...,0)$, where the only nonzero value is $a^{(k)}(Z_{t})$ at the $2k$-coordinate. Recall that we are working under the hypothesis that $a(Z_{t}^{k},\Pi_{t}^{N})$ can not be degenerated, which implies that $a^{(k)}(Z_{t})$ is lower bounded away from 0, for any $k\in\{1,...,N\}$.
This notation is convenient because, to demonstrate the hypoellipticity of our system, we will impose some nondegeneracy conditions on the coefficients in (10). These conditions will ensure that the system is hypoelliptic according to Definition 1 below. To introduce these constraints, consider the following vector fields on $\mathbb{R}^{2N}$ associated with the coefficients of Equation (10):
$\begin{cases}A_{j}:=\sum_{k=1}^{2N}A_{j}^{k}(z)\partial_{z_{k}},\quad j=1,...,2N\\ B:=\sum_{k=1}^{2N}B^{k}(z)\partial_{z_{k}}.\end{cases}$
@@
$A_{j}^{\nabla}A_{h}:=\sum_{l,k=1}^{2N}A_{j}^{l}\partial_{l}A_{h}^{k}\partial_{z_{k}}.$ (11)
@@
We will now use the Lie brackets to express the stochastic differential equation (10) in terms of the Stratonovich integral instead of the Ito integral. For further details about the Stratonovich integral, refer to [19, Section 7] and [81]. Given two real-valued continuous semimartingales $X$ and $Y$, we denote the Stratonovich differential (see [81, Definition 1.2.8]) as:
$X\circ\mathrm{d}Y:=X\mathrm{d}Y+\frac{1}{2}\mathrm{d}[X,Y],$ (12)
where $\mathrm{d}X$ is the standard Ito differential and $[X,Y]$ is the quadratic covariation.
With the Stratonovich notation, the Ito formula can be written as a standard chain rule (see [81, Proposition 1.2.10]). With this notation in mind, let us introduce the vector field $A_{0}$, which appears when we write the model in (10) in terms of the Stratonovich integral:
$A_{0}:=\sum_{k=1}^{2N}\left(B^{k}(z)-\frac{1}{2}\sum_{j,l=1}^{2N}A_{l}^{j}(z)\partial_{j}A_{l}^{k}(z)\right)\partial_{z_{k}}=B-\frac{1}{2}\sum_{l=1}^{2N}A_{l}^{\nabla}A_{l}.$
It is straightforward to check that replacing the Itô integral in (10) with the Stratonovich integral yields:
$dZ_{t}=A_{0}(Z_{t})dt+\sum_{j=1}^{2N}A_{j}(Z_{t})\circ dB_{t}^{j}.$
At this point, we are finally ready to define hypoellipticity in the sense of the stochastic calculus of variations (see p.129 of [68]).
@@
A differential operator $\mathcal{L}$ on an open set $G$ of $\mathbb{R}^{m}$ with smooth coefficients is hypoelliptic if, whenever $\nu$ is a distribution on $G$, $\nu$ is a smooth function on any open set $G^{\prime}\subset G$ on which $\mathcal{L}u$ is smooth.
@@
Therefore, it suffices to prove that we can apply Hörmander’s theorem to achieve the desired hypoellipticity of our model. This holds under a non-degeneracy hypothesis on the derivatives of the drift coefficients.
@@
$\forall(y,x)\in\mathbb{R}^{N}\times\mathbb{R}^{N}$, $\forall k,i\in\{1,...,N\}$, $\partial_{x_{k}}b_{1}^{(i)}(y,x)\neq 0$.
Remark 4.1.
One might wonder whether, instead of studying the hypoellipticity of the full system, it would be possible to examine the hypoellipticity of each particle separately (i.e., the two-dimensional SDE). This approach can be pursued when focusing on the limiting McKean–Vlasov system, for which the same argument as in Section 2.1 of [75] applies; in this setting, the dependence of the coefficients on the law does not introduce additional complications. For interacting particles, however, the situation is more delicate: the coefficients depend on the empirical measure, mixing the randomness of different particles, and making it unclear how to write the second-order differential operator for a single particle. For this reason, we investigate the hypoellipticity of the $2N$-dimensional SDE rather than that of each particle individually.
We also note that Hörmander’s condition in high dimensions has been established in related contexts (see, in particular, p. 366 of [26]). However, the models differ: in [26] the rough component has dimension $p$ and the smooth component dimension 1, whereas here both components of each particle have dimension 1, yielding a $2N$-dimensional system where rough and smooth components have equal weight in the basis construction. Thus, in [26] one builds a basis for $\mathbb{R}^{p+1}$, while in our case a basis for $\mathbb{R}^{2N}$ is required. Although the proof ideas are similar, the difference in structure leads to technical distinctions (reflected, for instance, in the fact that our Hypothesis A8 does not coincide with Condition 1 in [26]). We therefore state our result formally in Proposition 4.2, whose proof is given in Section 7.
@@
Assume that A8 holds true. Then the vectors
$(A_{2},[A_{0},A_{2}],A_{4},[A_{0},A_{4}],...,A_{2N},[A_{0},A_{2N}])=(A_{2k},[A_{0},A_{2k}])_{k\in\{1,...,N\}}$
form a basis for $\mathbb{R}^{2N}$ and the system is hypoelliptic.
The discussion above is crucial for reducing the model (9) to an integrated diffusion. Specifically, one can apply the change of variable $\hat{X}_{t}^{(i)}=b_{1}^{(i)}(Y_{t},X_{t})$. Consequently, for any $i\in\{1,\ldots,N\}$, the first equation in (9) becomes:

### v3

@@ 3 Estimators and main results
The condition $N\Delta_{n}\to 0$ is essential for the asymptotic normality in Theorem 3.2, serving to approximate the contrast derivative with martingale increments and to manage inter-particle correlations. In the classical setting where $T=n\Delta_{n}\to\infty$, such constraints (e.g., $n\Delta_{n}^{2}\to 0$) have been progressively relaxed using higher-order approximations [31, 87, 56, 51], with similar developments for jump diffusions [5, 6, 38, 80] and interacting particle systems [4]. However, while higher-order schemes might mitigate discretization errors, they do not resolve the intrinsic particle correlations. In our partially observed framework, the dependency between $X$ and $\tilde{X}$ further complicates this relaxation, which we leave for future investigation.
Our primary goal is to initiate statistical inference for kinetic interacting particle systems, a field currently less mature than its non-interacting counterpart. In this work, we employ a locally degenerate approximation for the contrast function, based on the Euler–Maruyama discretization of the intractable process.
Alternative strategies in the non-interacting setting (see e.g., [26, 39, 49, 65, 73]) utilize locally non-degenerate approximations via transition density schemes. While such approaches accommodate broader hypoelliptic models, their asymptotic analysis for interacting particle systems is significantly more complex. To maintain a focused scope, we prioritize the locally degenerate framework. Notably, the very recent contribution [52], appearing during our revision, extends these ideas to the non-degenerate case. We view this as evidence of the timeliness of our work and its role in stimulating further research into broader classes of degenerate interacting particle systems.
4 Motivating model examples
In this section, we provide several canonical examples from statistical physics, engineering, and econophysics to demonstrate the versatility of the model framework introduced in (3).
4.1 Interacting underdamped Langevin dynamics
The primary physical realization of our framework is the interacting underdamped Langevin dynamics. This model describes $N$ particles subject to a confining potential $V$, a reciprocal pair-interaction potential $W$, and a dissipative friction force. The drift and diffusion coefficients are given by:
$b(Z^{i}_{t},\Pi^{N}_{t})$ $=-\gamma X^{i}_{t}-V^{\prime}(Y^{i}_{t})-\frac{1}{N}\sum_{j=1}^{N}W^{\prime}(Y^{i}_{t}-Y^{j}_{t}),$ (9)
$a(Z^{i}_{t},\Pi^{N}_{t})$ $=\sqrt{2\gamma\beta^{-1}},$
where $\gamma>0$ is the friction coefficient and $\beta^{-1}$ represents the temperature of the environment. The drift term $b$ encapsulates the competition between conservative forces and non-conservative dissipation, serving as a microscopic foundation for Vlasov-type kinetic equations.
4.2 The kinetic mean-field Ornstein–Uhlenbeck process
A fundamental benchmark for analytical tractability and hypocoercivity theory is the “harmonic oscillator” of kinetic systems. By assuming quadratic potentials of the form $V(y)=\frac{\kappa}{2}y^{2}$ and $W(y)=\frac{\alpha}{2}y^{2}$, the drift becomes linear in the state variables:
$dX^{i}_{t}=\left(-\gamma X^{i}_{t}-\kappa Y^{i}_{t}-\alpha(Y^{i}_{t}-\bar{Y}^{N}_{t})\right)dt+\sqrt{2\gamma\beta^{-1}}dB^{i}_{t},$ (10)
where $\bar{Y}^{N}_{t}=\frac{1}{N}\sum_{j=1}^{N}Y^{j}_{t}$ represents the center of mass. Notice that $\bar{Y}^{N}_{t}$ is a direct function of the empirical measure $\Pi^{N}_{t}$, given by $\bar{Y}^{N}_{t}=\int_{\mathbb{R}\times\mathbb{R}}y\,\Pi^{N}_{t}(dx,dy)$. This model is unique as it preserves the Gaussianity of the distribution, providing a framework for studying convergence rates to equilibrium in degenerate diffusion settings.
4.3 Inertial Kuramoto models in power systems
Beyond statistical mechanics, the framework naturally extends to synchronization problems in power systems engineering. The synchronization of electrical generators can be mapped to our model by interpreting $Y^{i}$ as the phase angle and $X^{i}$ as the frequency deviation. Near the synchronized manifold, the sinusoidal coupling linearizes to:
$b(Z^{i}_{t},\Pi^{N}_{t})=-\gamma X^{i}_{t}-K(Y^{i}_{t}-\bar{Y}^{N}_{t}).$ (11)
This specific formulation is essential for assessing the resilience and stability of power grids against stochastic fluctuations in demand or renewable energy supply.
4.4 Systemic risk in interbank networks
Finally, a particular case of (10) is highly relevant in econophysics for modeling the stability of interbank reserves. Let $Y^{i}_{t}$ denote the log-monetary reserves of a given bank and $X^{i}_{t}$ its lending rate. Regulatory pressures and market averages act as harmonic restoring forces:
$b(Z^{i}_{t},\Pi^{N}_{t})=-\gamma X^{i}_{t}-\alpha(Y^{i}_{t}-\bar{Y}^{N}_{t}).$ (12)
Here, the “harmonic” attraction toward the market average $\bar{Y}^{N}_{t}$ represents a mean-reverting mechanism that prevents individual bank insolvency, while the noise $dB^{i}_{t}$ accounts for idiosyncratic market shocks.
@@ 5 On the Hormander’s condition for hypoellipticity
First, let us introduce the empirical projection of a function defined on measures. For any real-valued function $\phi:\mathcal{P}_{2}(\mathbb{R}^{l})\to\mathbb{R}$ (with $l\in\mathbb{N}$), we define its lifted version $\phi^{N}:\mathbb{R}^{lN}\to\mathbb{R}$ acting on the particle system by:
$\phi^{N}(z_{1},\ldots,z_{N}):=\phi\left(\frac{1}{N}\sum_{k=1}^{N}\delta_{z_{k}}\right).$
This definition allows us to view the coefficients of the McKean–Vlasov equation as functions of the particle configuration $Z_{t}=(Z_{t}^{1},\dots,Z_{t}^{N})$.
Specifically, for the drift coefficient $b_{j}(x,\mu)$ (where $j=1,2$), we define the corresponding function on the particle system, denoted by $b_{j}^{(i)}:\mathbb{R}^{2N}\to\mathbb{R}$, as:
$b_{j}^{(i)}(Z_{t}):=b_{j}\left(Z_{t}^{i},\frac{1}{N}\sum_{k=1}^{N}\delta_{Z_{t}^{k}}\right).$
Here, the dependency on the measure argument is replaced by the empirical measure of the system $Z_{t}$. Similarly, the diffusion coefficient $a(x,\mu)$ induces a function $a^{(i)}:\mathbb{R}^{2N}\to\mathbb{R}$ defined by:
$a^{(i)}(Z_{t}):=a\left(Z_{t}^{i},\frac{1}{N}\sum_{k=1}^{N}\delta_{Z_{t}^{k}}\right).$
@@
$dZ_{t}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{B}}(Z_{t})dt+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}(Z_{t})dW_{t},$ (14)
where $W_{t}$ is a $2N$-dimensional standard Brownian motion. Here,
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{B}}(Z_{t})=(b_{1}^{(1)}(Z_{t}),b_{2}^{(1)}(Z_{t}),...,b_{1}^{(N)}(Z_{t}),b_{2}^{(N)}(Z_{t}))\in\mathbb{R}^{2N},$
[…]
This notation is convenient because, to demonstrate the hypoellipticity of our system, we will impose non-degeneracy conditions on the coefficients in (14). These conditions will ensure the system is hypoelliptic according to Definition 1 below. To introduce these constraints, consider the following vector fields on $\mathbb{R}^{2N}$ associated with the coefficients of Equation (14):
$\begin{cases}A_{j}:=\sum_{k=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{j}^{k}(z)\partial_{z_{k}},\quad j=1,...,2N,\\ B:=\sum_{k=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{B}}^{k}(z)\partial_{z_{k}}.\end{cases}$
@@
$A_{j}^{\nabla}A_{h}:=\sum_{l,k=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{j}^{l}\partial_{l}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{h}^{k}\partial_{z_{k}}.$ (15)
@@
We will now use these vector fields to express the stochastic differential equation (14) in terms of the Stratonovich integral instead of the Itô integral. For further details about the Stratonovich integral, we refer to [19, Section 7] and [83]. Given two real-valued continuous semimartingales $X$ and $Y$, the Stratonovich differential (see [83, Definition 1.2.8]) is given by:
$X\circ dY:=XdY+\frac{1}{2}d[X,Y],$ (16)
where $dX$ is the standard Itô differential and $[X,Y]$ is the quadratic covariation.
Under the Stratonovich formulation, the Itô formula simplifies to the standard chain rule (see [83, Proposition 1.2.10]). With this in mind, let us introduce the vector field $A_{0}$, which emerges from the Itô-to-Stratonovich conversion of the model in (14):
$A_{0}:=\sum_{k=1}^{2N}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{B}}^{k}(z)-\frac{1}{2}\sum_{j,l=1}^{2N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{l}^{j}(z)\partial_{j}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{A}}_{l}^{k}(z)\right)\partial_{z_{k}}=B-\frac{1}{2}\sum_{l=1}^{2N}A_{l}^{\nabla}A_{l}.$ (17)
It is straightforward to verify that rewriting (14) as a Stratonovich equation yields:
$dZ_{t}=A_{0}(Z_{t})dt+\sum_{j=1}^{2N}A_{j}(Z_{t})\circ dW_{t}^{j}.$ (18)
At this point, we are equipped to define hypoellipticity in the context of the stochastic calculus of variations (see [70, p. 129]).
@@
A differential operator $\mathcal{L}$ on an open set $G\subset\mathbb{R}^{m}$ with smooth coefficients is hypoelliptic if, for any distribution $\nu$ on $G$, $\nu$ is a smooth function on any open subset $G^{\prime}\subset G$ where $\mathcal{L}\nu$ is smooth.
@@
Therefore, achieving the desired hypoellipticity of our model reduces to applying Hörmander’s theorem. This holds under a non-degeneracy hypothesis on the derivatives of the drift coefficients.
@@
For all $(y,x)\in\mathbb{R}^{N}\times\mathbb{R}^{N}$ and for all $k,i\in\{1,...,N\}$, $\partial_{x_{k}}b_{1}^{(i)}(y,x)\neq 0$.
Although analyzing the hypoellipticity of individual particles as isolated two-dimensional SDEs is valid for the limiting McKean–Vlasov system ([77, Section 2.1]), this decoupled approach is unsuitable for interacting particles. Because the coefficients depend on the empirical measure, the noise across different particles is coupled, precluding a standard definition of the second-order differential operator for an isolated particle. Consequently, we investigate the hypoellipticity of the joint $2N$-dimensional SDE.
We also note that Hörmander’s condition in high dimensions has been established in related contexts. We refer to [26, p. 366] and to the insightful discussions in [50, 51, 52]. However, our model differs structurally: in [26], the rough component has dimension $p$ and the smooth component has dimension $1$, whereas here, both components of each particle have dimension $1$. This yields a $2N$-dimensional system where the rough and smooth components carry equal weight in the basis construction. Thus, while [26] builds a basis for $\mathbb{R}^{p+1}$, we require a basis for $\mathbb{R}^{2N}$. Although the underlying ideas are similar, this structural difference leads to technical distinctions (e.g., our Hypothesis A8 does not coincide with Condition 1 in [26]). We formalize our result in Proposition 5.1, with the proof deferred to Section A.
@@
Assume that A8 holds true. Then, at every point $z\in\mathbb{R}^{2N}$, the vectors associated with the vector fields
$(A_{2},[A_{0},A_{2}],A_{4},[A_{0},A_{4}],\dots,A_{2N},[A_{0},A_{2N}])=(A_{2k},[A_{0},A_{2k}])_{k\in\{1,\dots,N\}}$
span the tangent space $T_{z}\mathbb{R}^{2N}\cong\mathbb{R}^{2N}$. Consequently, Hörmander’s condition is satisfied and the system is hypoelliptic.
The algebraic condition established in Proposition 5.1 carries a clear probabilistic intuition that is well-known in the study of degenerate diffusions (see, e.g., [49]). The fact that the vector fields $A_{2k}$ and their Lie brackets $A_{0},A_{2k}]$ span $\mathbb{R}^{2N}$ reflects how the stochastic driving forces propagate through the entire system. Specifically, the basis vectors $A_{2k}$ correspond to the direct Brownian noise $W^{k}_{t}$ acting on the velocity coordinates $X^{k}_{t}$. The Lie brackets $A_{0},A_{2k}]$, on the other hand, capture the integrated noise $\int_{0}^{t}W^{k}_{s}ds$. This integrated noise propagates into the position coordinates $Y^{k}_{t}$ through the kinematic relation $dY^{k}_{t}=X^{k}_{t}dt$, as can be readily seen from an Itô–Taylor expansion. Consequently, the noise effectively regularizes all $2N$ coordinates of the phase space, leading to the existence of a smooth Lebesgue density for the $2N$-dimensional system.
The proposition above is crucial for reducing the model (13) to an integrated diffusion. Specifically, one can apply the change of variable $\hat{X}_{t}^{(i)}=b_{1}^{(i)}(Y_{t},X_{t})$. Consequently, for any $i\in\{1,\ldots,N\}$, the first equation in (13) becomes: